OpenStudy (anonymous):

Following equation (see Reply), Is there an better method work or process to figuring it out? This has a solution, But I am looking for better&simpler methodologies and processes to get the solution. Keep answer as an irrational Fraction

OpenStudy (anonymous):

\[\sqrt{\frac{ 11 }{ 2 } - 3\sqrt{2}}\] \[\sqrt{ (\frac{ 2 }{ 2 } + \frac{ 9 }{ 2 }) - 3\sqrt{2} } = \sqrt{ (\frac{ 2 }{ 2 } + \frac{ 9 }{ 2 }) - \frac{6}{2}\sqrt{2} } \] \[\sqrt{ (\frac{ 2 }{ 2 } + \frac{ 9 }{ 2 }) - \frac{3}{2}\sqrt{2} - \frac{3}{2}\sqrt{2} } = \sqrt{\frac{2}{2} + ((-)\frac{3}{2})^{2} - \frac{3}{2}\sqrt{2} - \frac{3}{2}\sqrt{2}} \] Then \[\sqrt{\frac{ 11 }{ 2 } - 3\sqrt{2}}\ =1+ \frac{ 3 }{ 2 }\sqrt{2}\] Im having issues getting the middle part and figuring if there is an easier method to figuring this out.

OpenStudy (jiteshmeghwal9):



OpenStudy (jiteshmeghwal9):

\[\LARGE{\sqrt{5.5-3\sqrt{2}}}\]

OpenStudy (jiteshmeghwal9):

\[\sqrt{5.5-3(1.414)}\]

OpenStudy (jiteshmeghwal9):

not sure but i think it can make easy the solution :)

OpenStudy (anonymous):

Nah, as I said in the equation \[\sqrt{\frac{ 11 }{ 2 } - 3 \sqrt{2}} = 1 - \frac{ 3 }{2 }\sqrt{2}\] thus the answer is \[1 - \frac{ 3 }{2 }\sqrt{2}\] square root cancels out the exponential \[\sqrt{ (1 - \frac{ 3 }{2 }\sqrt{2})^2}\] The answer is to be given as an irrational fraction to simplify and give more accurate result than having it in decimal . What I am trying to better my self is in the process of solving this equation to its simplest form:

OpenStudy (anonymous):

Also I accidentally miswrote the solution to the equation in the reply, it should be the above.

OpenStudy (anonymous):

I guess my solution is as advance as people here can explain it. Will keep this open until I get a better solution to this